A car moves for half of its time at 80 km/h and for rest half of
time at 40 km/h. Total distance covered by car is 60 km. What
is the average speed of the car ?
(a) 60 km / h
(b) 80 km / h
(c) 120 km / h
(d) 180 km / h
 
What is the correct answer? 

Grade:12th pass

5 Answers

g v subba reddy
7 years ago
time taken to cover half distance with speed 80=distance /time
(1/2*60)/80=3/8hours
time taken to cover half distance with speed 40=30/40=3/4
average speed =total distance /time
=60/(3/8+3/4)=53.3
Venky
7 years ago
if a body travels half time with velocity v1& another half time with v2 then ava. velocity is (v1+v2)/2 so (80+40)/2=60
CHALLA NAVNEESWAR REDDY
7 years ago
given v1=80km/h, v2=40km/h, distance=60km
we know that
average speed=(v1+v2)/2
=(80+40)/2, =120/2,  =60km/h
Zayn Sheikh
6 years ago
As We know that Distance = Speed × Time So, here D》》 80T+40T = 60 Hence T= 0.5 So , total time = 2× 0.5 which is 1 secTherefore Average speed = 60/1 = 60 km/hr
nimra
5 years ago
IN FIRST HALF                          S=V1T1              60=80T/2   =40T
IN SECOND HALF                     S=V2T2             60=40T/2     =20T
                          S =V1T1 +V2T2 
                          60=60T 
                         T=IHOUR
                         V=TOTAL DISTANCE /TOTAL TIME 
                          60 KM/HOUR